If Rs 750 amount to Rs 885 in three years at simple interest, what will Rs 1200 amount to in 3 and half years (The rate is the same in both cases.)
Answer by rfer(16322)

(885-750)=750r3
135=2250r
135/2250=r
r=0.06
--------------
I=1200*0.06*3.5
I=252
----------------
1200+252=1452